A certain fluid in a closed system initially at a pressure of 200 psia and having a volume of 3 ft3 undergoes a frictionless process during which the product pV1.3 = C. If p2 = 40 psia, compute the work done by or on the system.


1
Expert's answer
2021-01-21T03:52:30-0500

pV1.3 = c

p1 = 200 psia = 1.38 x 106 Pa

V1 = 3 ft3 = 0.085 m3

p2 = 40 psia = 2.76 x 105 Pa

W = "\smallint" pdV = p1V1n"\smallint" dV/Vn = (p2V2 - p1V1) / (1 - n) = (p1V1 - p2V2) / (n - 1)

c = p1V11.3 = 1.38 x 106 Pa x (0.085 m3)1.3 = 847.5

V2 = (c/p2)1/1.3 = (847.5 / 2.76 x 105)1/1.3 = 0.011 m3

W = (1.38 x 106 x 0.085 - 2.76 x 105 x 0.011) / (1.3 - 1) = 380 880 J = 381 kJ